Solve the system using the addition method. If there is no solution or if there are many solutions and the system's equations are dependent, so state. The three equations are:
6x-y+3z+22
x+3y-z=-4
3x+3y-4z=-9
I am having to do my homework on course compass (AKA mathlab) and only get 3 chances to enter in my answer before it giving me a new problem. This is what I have done so far to get an answer and it telling me my answer is wrong.
I took equation 1 and 3 and eliminated X and got another equation which is now equation 4. I then took equation 2 and 3 and eliminated X and got my 5th equation. I then added equation 4 and 5 and eliminated Z which gave my my Y answer. I then plugged my Y answer in to equation 5 and got my Z. I then took my Y and Z answers and plugged them into equation 1 to get my X. Somewhere in the mess of it all, I got the wrong answers. Can you help me?
Thanks,
Lexi

Answer by scott8148(6628)

4___ -7y + 11z = 40

5___ -6y - z = 3

-73y = 73 ___ y = -1

-6(-1) - z = 3 ___ z = 3

x + 3(-1) - (3) = -4 ___ x = 2